Find two consecutive even integers whose product is 48.
and

Find two consecutive odd integers whose product is 63.

Answer by Alan3354(69443)

Find two consecutive even integers whose product is 48.
-----------------
The quick and easy way:
sqrt(48) is ~7 (7*7=49).
It's the numbers above and below 7,
6 and 8 (and -8 and -6)
----------
The hard way:
n*(n+2) = 48
n^2 + 2n - 48 = 0
(n+8)*(n-6) = 0
--> 6 & 8 and -8 & -6
------------
and
Find two consecutive odd integers whose product is 63.
Do this one the same way.